OpenStudy (grandmoffchris):

What is -5/6 ÷ 9/10 ? A. -27/25 B. -4/3 C. -3/4 D. -25/27

OpenStudy (grandmoffchris):

Will fan and medal pls

OpenStudy (grandmoffchris):

@dan815 @Abstract_Undercover

OpenStudy (anonymous):

Wht do u think it is

OpenStudy (grandmoffchris):

Well -3/4

OpenStudy (grandmoffchris):

Is my answer but im not sure

OpenStudy (grandmoffchris):

Help!

OpenStudy (mathstudent55):

To divide fractions, change division into multiplications, and flip the second fraction. Then multiply the fractions.

OpenStudy (grandmoffchris):

im horrible with fractions

OpenStudy (grandmoffchris):

dividing em, multipling em, im bad at em

OpenStudy (mathstudent55):

Example: \(- \dfrac{2}{3} \div \dfrac{7}{11} = -\dfrac{2}{3} \times \dfrac{11}{7} = -\dfrac{22}{21}\)

OpenStudy (grandmoffchris):

Have no idea

OpenStudy (anonymous):

solve it @GrandMoffChris

OpenStudy (grandmoffchris):

Well Ok

OpenStudy (grandmoffchris):

i dont get it

OpenStudy (grandmoffchris):

i really dont

OpenStudy (grandmoffchris):

Nvm, got it wrong

OpenStudy (grandmoffchris):

u can leave this ok?

OpenStudy (mathstudent55):

\(-\dfrac{5}{6} \div \dfrac{9}{10} \) We need to divide the first fraction by the second fraction. Step 1. We copy the first fraction as it is. \(-\dfrac{5}{6} \) Step 2. We switch the division into a multiplication. \(-\dfrac{5}{6} \times\) Step 3. We flip the second fraction. \(-\dfrac{5}{6} \times \dfrac{10}{9} \) Step 4. We multiply the fractions. \(-\dfrac{5}{6} \times \dfrac{10}{9} = - \dfrac{5 \times 10}{6 \times 9} = -\dfrac{50}{54} \) The multiplication ios done. Now we see if we can reduce the fraction. Step 5. Try to reduce the result. We notice that both 50 and 54 are divisible by 2. \(-\dfrac{50}{54} = -\dfrac{25 \times 2}{27 \times 2} = -\dfrac{25}{27}\)